### v2

# Enhanced Optimization Strategies to Design an Underactuated Hand Exoskeleton

Abstract
Exoskeletons can boost human strength and provide assistance to individuals with physical disabilities. However, ensuring safety and optimal performance in their design poses substantial challenges. This study presents the design process for an underactuated hand exoskeleton, first including a single objective (maximizing force transmission), then expanding into multi-objective (also minimizing torque variance and actuator displacement). The optimization relies on a Genetic Algorithm, the Big Bang-Big Crunch Algorithm, and their versions for multi-objective optimization. Analyses revealed that using Big Bang-Big Crunch provides high and more consistent results in terms of optimality with lower convergence time. In addition, adding more objectives offers a variety of trade-off solutions to the designers, who might later set priorities for the objectives without repeating the process – at the cost of complicating the optimization algorithm and computational burden. These findings underline the importance of performing proper optimization while designing exoskeletons, as well as providing a significant improvement to this specific robotic design.
keywords:
Hand Exoskeleton , Rehabilitation Robotics , Evolutionary Computation , Multi-Objective Optimization , Design Optimization
journal: ArXiv Preprint
[…]
1 Introduction
Exoskeletons are wearable robotic devices that can improve users’ functional abilities during physically challenging tasks [poliero2020applicability, butzer2021fully] or assist/rehabilitate them in case of functional disabilities [stroppa2017robot, huang2015design, pons2010rehabilitation]. They are usually designed to be worn on a specific body location; thus, their proper design is crucial for force transmission efficiency but also, and most importantly, for users’ safety [sarac2019design]. For an exoskeleton to be safe, the anatomic and mechanical joints must be aligned during possible (intentional or unintentional) movements of users [van2015considerations, fisahn2016effectiveness]. In addition, the exoskeleton design must be optimized to provide the highest level of well-distributed output forces through small and lightweight actuators.
The engineering design is usually performed through mathematical optimization, searching for the best element (optimum) within a set of alternatives (decision variables’ values) based on one or more specific criteria (objectives) [sioshansi2017optimization, statnikov2012multicriteria, andersson2000survey]. Traditional optimization methods rely on numerical and calculus-based techniques [bonnans2006numerical] and might be challenging for exoskeleton design [stroppa2023optimizing] due to properties such as non-discrete domains, multimodality, discontinuity [norde2000characterizing], and the need for reliability [dizangian2015reliability] and robustness [nomaguchi2016robust]. Alternatively, the nature-inspired Evolutionary Computation (EC) methods – particularly Evolutionary Algorithms (EAs) – are well-known and effective in dealing with engineering optimization problems [dumitrescu2000evolutionary, wang2017autonomous] and often with exoskeleton design [stroppa2023optimizing, datta2015analysis] and other robotic devices [stroppa2024optimizing].
Figure 1: The optimized device for the index finger component of U-HEx.
Previously, we proposed optimizing an underactuated hand exoskeleton (U-HEx [sarac2017design], shown in Fig. 1) using EAs, in which we retrieved the best set of link lengths (decision variables) that maximize the output force transmission (objective function) [akbas2024impact]. We specifically chose EAs over other nonlinear numerical methods due to their robustness in handling black-box optimization and their ability to efficiently explore the design space of our underactuated hand exoskeleton without requiring explicit gradients or analytical formulations for the objective functions – especially considering that the objective functions come from a motion simulation of the kinematic chain. While the first design of U-HEx was optimized with a naive-deterministic search approach, which iteratively analyzed each possible combination of design parameters (i.e., “brute force”), we employed a Genetic Algorithm (GA) [goldberg1987genetic, goldberg1990real] and the Big Bang-Big Crunch algorithm (BBBC) [erol2006new]. Comparative analyses revealed that both EAs yielded more consistent and optimal solutions than brute force in a significantly shorter time. This allowed us to extend the search space by including more decision variables (i.e., three more link lengths), significantly improving U-HEx’s efficiency.
Yet, our previous work had a few limitations. Firstly, we only focused on a single objective function. Secondly, we did not constrain the desired actuator movements to reach a natural range of motion for the user’s hand. In other words, we have found an optimal set of link lengths with no prior assumption on the desired actuator displacements. This assumption cannot be overlooked, since our design involves a linear actuator with a physical limit to how much it could move from its base – so that it can fit on top of the hand [sarac2017design]. Therefore, it is crucial to minimize the desired range of actuator displacement to assist natural ranges of motion for the finger joints without sacrificing the force transmission performance. On the other hand, we have previously constrained the ratio between the actuated finger joints in our optimization problem, while it would be preferred to be as close to 1 for well-distributed and safe force transmission.
In this study, we have built up on our Single-Objective Optimization Problem (SOOP) [akbas2024impact] by improving U-HEx’s design in two alternative design processes:
  1. 1.
    we redefined the SOOP with an additional constraint to the actuator displacement based on the original design [sarac2017design], and we hypothesize that this would result in a lower overall transmission and potentially different trends between optimization methods – even though we have very effective optimization methods in hand ($H_{1}$); and
  2. 2.
    we extended the SOOP to a Multi-Objective Optimization Problem (MOOP) by also optimizing the balance between the torques applied to each finger joint and minimizing the required actuator displacement (i.e., turning constraints into objectives), and we hypothesize that treating constraints as further objectives would better accommodate the problem and lead to a trade-off set of different designs balancing the device’s properties ($H_{2}$).
Our contributions involve (i) providing a set of optimized designs for a state-of-the-art exoskeleton that was originally optimized with inefficient methods [sarac2017design], (ii) providing a walkthrough for a robotic design process that can be used as a guideline by future and less experienced robotic designers, (iii) comparing different algorithms and strategies for MOOPs to determine which one is most suitable for our design problem, and (iv) introducing two novel EC methods (NS-BBBC and SP-BBBC) suitable for robotic optimization, which are based on a state-of-the-art algorithm for single-objective optimization (i.e., BBBC). Table 1 summarizes the different contributions on each work of the U-HEx device, comparing the current work with its state-of-the-art.
Table 1: Comparison of different works novelties on U-HEx
Reference Novelty Optimization Method
Sarac et al. 2017 [sarac2017design] Kinematics and mechanical design Brute Force
Akbas et al. 2024 [akbas2024impact] Optimized device based on maximum torque magnitude Brute Force, GA, BBBC
Current Work Comprehensive and broader optimization of the device on multiple criteria GA, BBBC, NSGA-II, SPEA2, NS-BBBC, SP-BBBC
The rest of the work is organized as follows: Sec. 2 defines the optimization design and provides a background on the methods we used; Sec. 3 describes the first optimization problem with a single objective (i.e., maximize the force transmission); Sec. 4 describes the second optimization problem with two more objectives (i.e., minimize the distribution of forces and actuator displacement); Sec. LABEL:sec:comparison discusses the differences in the retrieved designs and the impact of each optimization process; and Sec. LABEL:sec:conclusion summarizes the work and presents future ideas.
2 Exoskeleton Design Optimization
2.1 U-HEx Kinematics
U-HEx in Fig. 1 is a hand exoskeleton designed for rehabilitation [sarac2017design]. It was designed as a linkage structure with mechanical joints intentionally misaligned from the anatomical joints of the user. This design choice has multiple advantages. Firstly, the misalignment offers a much safer operation, since the mechanical device does not impose hard forces on the joints directly. Any sort of undesired or extreme force would be transmitted to the fingers in a more natural way without the risk of harming the user. Secondly, thanks to this misalignment, U-HEx can adapt its operation automatically to users with different finger sizes as well. Finally, since U-HEx joints do not need to be aligned with the anatomical joints strictly, it is noticeably easy to wear.
The misalignments of U-HEx were designed carefully to achieve underactuation. A linear DC motor allows the device to close/open the finger in a direct manner. The idea behind underactuation is simple: the actuation is turned into motion at the first output joint until the first mechanical link gets in contact with an object. After the first link reaches an object and starts feeling the interaction forces, the spring-like joint coupling between output joints allows the mechanism to transmit the same actuation for the second output joint. Actuator forces are transmitted to output joints until the grasping is completed [Denizon2024, Yan2022]. From the point of hand exoskeletons, users’ fingers as a part of the mechanism: anatomical coupling between finger joints defines the required spring-like behavior for the underactuation [Li2023, Gu2024, sarac2017design]. Then, efficient force transmission requires the use of additional passive joints, which can assist motion for finger joints without aligning mechanical joints to the finger joints.
Fig. 2 represents the operation flow of U-HEx. The finger component is connected to the user’s finger and assists the finger in flexing until both finger joints are fully extended and the actuator stroke is at its minimum limit. In the beginning, as the actuator starts to move, the exoskeleton transmits forces to the proximal phalange and rotates the MCP joint. Encountering external forces along the proximal phalange caused by the physical interaction finalizes the rotation along the MCP joint. These forces transmit actuation forces to the intermediate phalange through the linkage-based mechanism with passive joints and rotating the PIP joint. Having two finger joints to control, the grasping task is completed when the PIP joint is rotated enough for the intermediate phalange to reach the object. Similarly, the extension of the finger starts by rotating the PIP joint first and the MCP joint later until the finger is totally extended and finger joints reach their physical limits. As a safety measure, the displacement stroke is designed to reach its minimum limit when both finger joints are extended to their natural limits.
Figure 2: Underactuation concept of U-HEx during a grasping task: (left) initial pose with all finger joints extended, (mid) actuator force moves the MCP joint until the first finger phalange gets in touch with the object, (right) actuator force is transmitted to move the PIP joint when the first finger phalange touches the object and finally grasping task is completed when both phalanges touch the object
Fig. 3(a) represents the kinematic structure of U-HEx. Each gray dot represents a passive rotational joint (or anatomical joint) while empty circles ($H$, $E$, and $C$) are fixed points along rigid links with $90^{o}$ fixed angle. The system has only one linear actuator between the points $O$ and $A$. The exoskeleton is fixed on the hand from points $K$ and $O$ with variable lengths in the $x$ and $y$ directions. Finally, the exoskeleton is attached to the finger phalanges from points $I$ and $J$ with passive linear joint sliders represented as $c_{1}$ and $c_{2}$, respectively.
Figure 3: Optimization process for U-HEx: for a given set of link lengths, a Simulink model is executed to compute numerical inverse kinematics and analytical statics as the finger joints are iterated from fully open to fully closed; while the optimization algorithms are implemented in MATLAB.
U-HEx design is optimized to retrieve the optimal link lengths following the schema and design process summarized in Fig. 3. The decision variables of the problem are the link lengths (depicted as red lines in the figure). The original U-HEx [sarac2017design] was optimized only for six link lengths: $\overline{BC}$, $\overline{CD}$, $\overline{DE}$, $\overline{EF}$, $\overline{FG}$, and $\overline{GH}$. due to the heavy computational requirements of brute force, which made it impossible to run the optimization with more link lengths. In our previous work [akbas2024impact], we included an additional three link lengths ($\overline{BK}$, $\overline{CI}$, and $\overline{EJ}$) in the optimization problem, potentially creating a more optimal solution. The bounds of the decision variables (i.e., the range of the link lengths) are reported in Table 2 and were chosen empirically [sarac2017design].
Table 2: Link Length Ranges / Decision Variable Bounds
Main Six Link Lengths (mm)
$\overline{BC}$ $\overline{CD}$ $\overline{DE}$ $\overline{EF}$ $\overline{FG}$ $\overline{GH}$
$38\rightarrow 60$ $10\rightarrow 30$ $15\rightarrow 51$ $15\rightarrow 51$ $27\rightarrow 56$ $64\rightarrow 100$
Additional Three Link Lengths (mm)
$\overline{BK}$ $\overline{CI}$ $\overline{EJ}$
$20\rightarrow 50$ $10\rightarrow 17$ $20\rightarrow 50$
Fig. 3 shows that at each step of the process, a new set of link lengths is assigned by the optimization method in MATLAB. This set is used to compute the numerical inverse kinematics [sarac2017design] as the finger joints are iterated from fully open ($0^{o}$) to fully closed ($80^{o}$ MCP; $90^{o}$ PIP) with a fixed-step solver. This process results in the corresponding movements around the joints (e.g., linear sliders $c_{1}$, $c_{2}$) and the desired actuator displacement ($L_{x}$). These movements are then used to compute the force transmission – i.e., expected torques around the finger joints ($\tau_{MCP},\tau_{PIP}$) for a unit actuator force (1 N) through analytical statics equations. Obtaining the force transmission triggers the MATLAB file to evaluate the obtained performance in terms of the constraints and the optimality. Depending on the performance, the optimization algorithm repeats this loop until the requirements are satisfied.
2.2 Optimization Algorithms (Evolutionary Algorithms)
In this paper, we will explore two optimization problems (SOOP and MOOP) using the same optimization algorithms based on EC. EC is a sub-field of soft computing offering nature-inspired algorithms: instead of updating on a single point in the search space, it generates a population of potential solutions (i.e., a set of different combinations of decision variables’ values) and evolves them toward the optimum using different metaheuristics (e.g., inspired by genetic recombination, natural selection, universe creation). Since the generic framework of EC methods does not rely on a specific mathematical model and can tackle objective functions defined implicitly through simulations, they are suitable for engineering problems such as optimizing U-HEx. Specifically, we will employ two different EAs, as described below.
2.2.1 Genetic Algorithm (GA)
GAs are the most widely used EAs because they directly replicate the process of natural selection and the concept of survival of the fittest [goldberg1989genetic, goldberg1990real]. As summarized in Algorithm 1, GA (i) creates a population of random solutions $P$ within the search space, (ii) assigns a fitness value to each solution based on the objective function, and (iii) produces new solutions $Q$ by combining the values of the solutions in the current population through a crossover process. Only the fittest solutions $M$ undergo crossover and are retained in subsequent generations, which allows the GA to evolve its population towards the optimal solution. Similar to biological processes, the crossover operation utilizes the characteristics of high-quality solutions (parents) to generate new, comparable solutions (offspring) – accelerating convergence to an optimum. However, there is no guarantee that the retrieved optimum is global rather than local. To address this issue, GAs incorporate an additional operator inspired by genetic mutation. This randomly alters the values of newly generated solutions, enhances search-space exploration, and helps in avoiding local optima.
input : Population size n, number of generations g
output : The most fitting solution $\textnormal{{P}}\left(1\right)$
1
2begin
3
[…]
6       for $i\leftarrow 1\text{ to }\textnormal{{g}}$ do
[…]
11
12
13      return P;
14
Algorithm 1 Genetic Algorithm
2.2.2 Big Bang-Big Crunch Algorithm (BBBC)
BBBC [erol2006new] is inspired by the universe’s evolution through two phases: explosion and implosion. The process involves energy dissipation, which creates disorder and randomness and then reorganizes this randomness into a new order. As summarized in Algorithm 2, BBBC (i) generates an initial random population $P$ uniformly distributed across the search space (the explosion, or “big bang”), and (ii) evaluates and collects the population into their center of mass $cm$ (the implosion, or “big crunch”). As the iterations progress, these two phases are repeated, with new solutions being generated closer to the center of mass, converging to the optimal solution. BBBC is recognized for its superior convergence speed compared to GAs, making it a suitable choice for our problem, given the high runtime of the objective function executed in a Simulink simulation.
input : Population size n, number of generations g
output : The most fitting solution $\textnormal{{P}}\left(1\right)$
1
2begin
3
[…]
5       for $i\leftarrow 1\text{ to }\textnormal{{g}}$ do
6             if $i\neq 1$ then
7                   $\textnormal{{P}}\leftarrow\textnormal{{bang}}\left(\textnormal{{cm}},i\right)$;
8
[…]
11
12
13      return P;
14
Algorithm 2 Big Bang-Big Crunch Algorithm
3 Design Process I: Single-Objective Optimization with Mechanical Constraints
U-HEx was designed with a linear actuator (Firgelli L12-P) with a maximum displacement of 50 mm [sarac2017design]. In this study, we evaluated the impact of presenting the mechanical limitations of the minimum and maximum linear actuator displacements on U-HEx design by comparing the optimality of their retrieved solutions against the original SOOP [akbas2024impact].
3.1 Optimization Problem (SOOP)
Similar to the previous work, the objective function is maximizing the force transmission during the operation:
maximize $\sqrt{\tau_{MCP}\text{ + }\tau_{PIP}}$ (1)
where $\tau_{MCP}$ and $\tau_{PIP}$ are calculated based on the values of the decision variables (link lengths). The optimization problem is subject to the following seven inequality constraints:
  • 1.
    $con_{1-4}$: U-HEx is connected to the user’s fingers through passive linear sliders ($c_{1}$ and $c_{2}$) whose movements should be limited by the finger size [sarac2017design], as Eqn. (2);
    $0\leq c_{1}\leq 35;\quad 0\leq c_{2}\leq 45$ (2)
  • 2.
    $con_{5-6}$: to guarantee a solid grasp and user’s safety, the torques $\tau_{MCP}$ and $\tau_{PIP}$ should have a ratio within a reasonable range [sarac2017design], as Eqn. (3); and
    $\frac{1}{15}\leq\frac{\tau_{MCP}}{\tau_{PIP}}\leq 15$ (3)
  • 3.
    $con_{7}$: to ensure a natural range of motion for the hand, the desired actuator displacement ($L_{x}$) must comply with the chosen actuator’s mechanical spectrum, as Eqn. (4) – this constraint was not considered in our previous work [akbas2024impact].
    $L_{x}\leq 50$ (4)
Each of these constraints is a function of the decision variables (link lengths) and is calculated in Simulink, similarly to the objective function (see Fig. 3(a) [sarac2017design]). Additional constraints to the problem are related to the decision variable bounds (i.e., the range of link lengths) reported in Table 2.
3.2 Experiment Outline (SOOP)
We designed an experiment with three factors as (i) the constraint to the desired actuator displacement (with and without $con_{7}$), (ii) the optimization algorithm (GA and BBBC), and (iii) the number of decision variables (six and nine link lengths). When the number of link lengths is limited to six, the additional three links are kept constant as $\overline{BK}=37$, $\overline{CI}=16$, and $\overline{EJ}=37$. These values were selected empirically as detailed in the original work [sarac2017design]. To ensure a valid comparison, we kept the algorithm parameters the same as our previous work [akbas2024impact], reported in Table 3. Each algorithm was executed 10 times on a computer with 16 core 5.4 GHz CPU and 64 GB RAM.
Table 3: Parameter Settings in SOOP for EAs
PARAMETER EXPERIMENT VALUE
Max N. of Generations (GA, BBBC) $50$
Population Size (GA, BBBC) $300$
Selection Type (GA) Binary Tournament [goldberg1991comparative]
Crossover Type (GA) blx-$\alpha$ ($\alpha$ = 0.5) [eshelman1993real]
Crossover Probability (GA) $1.0$
Mutation Type (GA) Polynomial [deb1996combined]
Mutation Probability (GA) $0.2$
Survival Strategy (GA) Elitist ($\mu+\lambda$ schema) [beyer2002evolution]
Crunch Method (BBBC) Best Fit [gencc2010big]
Constraint Handling (GA, BBBC) Deb’s Method [coello2002theoretical]
3.3 Evaluation Metrics (SOOP)
At each execution, we recorded the set of optimized link lengths (i.e., values of the decision variables belonging to the most fitting solution in the population) and compared each factor in terms of two evaluation metrics.
3.3.1 Optimality
In SOOP, optimality is the value of the objective function, defined in Eqn. (1).
3.3.2 Convergence Time (CT)
CT indicates the time needed for algorithms to find the optimal solution, defined in Eqn. (5).
$\text{CT}=\frac{\text{GC}}{\text{NG}}\cdot\text{RT}$ (5)
where the overall runtime (RT) indicates the overall time taken by the algorithm and the number of generations (NG) is 50 as indicated in Table 3. The generation of convergence (GC) is the observed generation in which the algorithm converged to a solution – specifically if the variance in the value of the retrieved solution did not change in the last 20 iterations with a margin of 0.05 Nm.
3.4 Results (SOOP)
We analyzed optimality and convergence time using a three-way Repeated Measure Analysis of Variance (RM-ANOVA). Fig. 4 shows the mean and standard error of the collected data, as reported in Table 4; whereas Table 5 summarizes the ANOVA results.
3.4.1 Results on Optimality
Fig. 4 shows that implementing the actuation constraint ($con_{7}$) significantly lowered the force transmission. The drop between with and without $con_{7}$ becomes higher with nine decision variables compared to six. We further analyzed the trends between optimization strategies and decision variables only with $con_{7}$ through a post-hoc analysis, as depicted by a zoomed-in version in Fig. 5. We found that the optimality results are statistically significantly higher with nine decision variables ($F(1,36)=1047621.937,p<0.001,\eta^{2}=1.000$) and while using BBBC ($F(1,36)=74.239,p<0.001,\eta^{2}=0.673$). In addition, we observed their interactions to be statistically significant ($F(1,36)=38.156,p<0.001,\eta^{2}=0.515$).
Figure 4: Force transmission (optimality) and convergence time comparisons between experiments (with a different number of variable link lengths): with the constraint to the maximum actuator displacement ($con_{7}=L_{x}\leq 50$) and without [akbas2024impact]. Plots report the mean and the standard error.
Table 4: Numerical Results of Design Process I (SOOP) with 6 and 9 Link Lengths (LL), with and without Actuation Displacement Constraint ($con_{7}$: $L_{x}\leq 50$)
Force Transmission (Nm)
without $con_{7}$ with $con_{7}$
GA (6 LL) $31.63\pm 0.01$ $28.39\pm 0.10$
BBBC (6 LL) $31.65\pm 0.05$ $28.52\pm 0.005$
GA (9 LL) $56.20\pm 0.41$ $28.61\pm 0.19$
BBBC (9 LL) $55.99\pm 0.54$ $29.43\pm 0.25$
Run Time (s)
without $con_{7}$ with $con_{7}$
GA (6 LL) $860.51\pm 168.25$ $1857.05\pm 32.84$
BBBC (6 LL) $312.50\pm 204.63$ $724.78\pm 306.87$
GA (9 LL) $1208.59\pm 78.31$ $1772.48\pm 182.43$
BBBC (9 LL) $802.50\pm 298.03$ $1391.63\pm 469.05$
3.4.2 Results on Convergence Time
Fig. 4 shows that adding $con_{7}$ to the optimization problem significantly increased the convergence time for both EAs and the number of decision variables. We further analyzed the data collected with $con_{7}$. We further analyzed the trends between optimization strategies and the number of link lengths only with $con_{7}$ through a post-hoc analysis. We found that the convergence time results are statistically significantly different with higher decision variables (($F(1,36)=9.001,p=0.005,\eta^{2}=0.200$)) and while using GA instead of BBBC ($F(1,36)=60.781,p<0.001,\eta^{2}=0.628$). In addition, we found their interactions to be statistically significant ($F(1,36)=14.989,p<0.001,\eta^{2}=0.294$).
Figure 5: A zoomed-in representation of force transmission (optimality) with the linear actuation constraint ($con_{7}$) as $L_{x}\leq 50$. The plots show the median, interquartile range, and outliers of the data.
3.5 Discussion (SOOP)
We investigated the impact of constraints based on the desired actuator displacement (limitations of the actuator chosen from the market, Firgelli L12-P) on the optimization problem using different EAs. We used two EAs (GA and BBBC) and compared our results to our previous study [akbas2024impact]. Our results (Fig. 4) show that including the actuation constraint lowers the optimality (i.e., force transmission) while using both optimization methods and decision variable settings (six and nine link lengths), as expected – therefore, $H_{1}$ holds true.
Fig. 4 also shows that increasing the number of decision variables significantly did not increase the optimality when the actuator constraint was introduced. This might indicate that all the solutions leading to high values of optimality require actuator displacements higher than 50 mm (i.e., unfeasible solutions), which will be investigated in the following design process with MOOP (see Sec. 4). In addition to a decrease in optimality, we observed that the actuation constraint resulted in statistically higher convergence times for the optimization algorithms. We speculate that higher convergence times might be due to the change in the search space and the computational toll of evaluating one more constraint [coello2002theoretical].
Table 5: Statistical Results of Design Process I (SOOP)
Optimality Convergence Time
Algorithm
(GA vs. BBBC)
F(1,72)=9.351,
p$=$0.003, $\eta^{2}$= 0.115
F(1,72)=45.624,
p$<$001, $\eta^{2}$= 0.388
N. of Link Lengths
(6 vs. 9)
F(1,72)=39221.829,
p$<$0.001, $\eta^{2}$= 0.998
F(1,72)=7.992,
p$=$0.006, $\eta^{2}$= 0.100
Actuator Constraint
(with vs. without)
F(1,72)=57325.641,
p$<$0.001, $\eta^{2}$= 0.999
F(1,72)=146.730,
p$<$0.001, $\eta^{2}$= 0.671
Algorithm $\times$ N. of Link Lengths
F(1,72)=3.209,
p$=$0.077, $\eta^{2}$= 0.043
F(1,72)=0.722,
p$=$0.398, $\eta^{2}$= 0.010
Algorithm $\times$ Constraint
F(1,72)=21.140,
p$<$0.001, $\eta^{2}$= 0.227
F(1,72)=0.164,
p$=$0.687, $\eta^{2}$= 0.002
N. of Link Lengths $\times$ Constraint
F(1,72)=35713.120,
p$<$0.001, $\eta^{2}$= 0.998
F(1,72)=11.982,
p$<$0.001, $\eta^{2}$= 0.143
Algorithm $\times$
N. of Link Lengths $\times$ Constraint
F(1,72)=13.668,
p$<$0.001, $\eta^{2}$= 0.160
F(1,72)=0.046,
p$=$0.830, $\eta^{2}$= 0.001
We also analyzed the differences between the two EAs and the number of decision variables when the actuation constraint was implemented separately since the optimality results were primarily influenced by the optimization problem without the actuation constraint. Fig. 5 shows that increasing the number of decision variables provides a statistical increase in force transmission – similar to our previous study [akbas2024impact].
Interestingly, we observed that BBBC offers statistically higher optimality results than GA while working with six and nine decision variables independently – even though previously we found no difference between them [akbas2024impact]. We speculate that this was caused by the GA being stuck in a local maxima instead of a global one. This is also supported by observing the convergence time: while BBBC reached the current optimal solution [^1] We remind readers that EC methods retrieve approximate optimal solutions rather than the exact one. relatively fast, GA continued the search for a significantly longer time. Thus, the convergence time of GA did not change as much as BBBC when the number of decision variables was increased. As mentioned in the state-of-the-art, BBBC was proposed to address some of the biggest disadvantages of GAs, such as premature convergence, convergence speed, and execution time [erol2006new].
4 Design Process II: Multi-Objective Optimization with Mechanical Constraints
The design process in SOOP (Sec. 3) showed the importance of considering the actuator displacement as a constraint. Introducing a constraint to the desired actuator displacement (Eqn. 4) reduced the force transmission of the exoskeleton but allowed it to be equipped with smaller actuators that can fit on top of the hand while still exerting a sufficient amount of force to assist users in opening and closing their fingers. In the meantime, users’ safety was ensured by balancing the torques applied on both finger joints, as detailed in Eqn. (3).
Even though these constraints allowed us to have a design within safe and feasible boundaries, they do not ensure reaching an optimal one about these factors. We re-defined the optimization problem to include two further objectives that were previously treated as constraints (i.e., $con_{5-6}$ and $con_{7}$). We believe that turning constraints into additional objectives can help find better solutions by allowing more flexibility and balancing multiple requirements effectively, resulting in a set of different designs with different properties.
4.1 Background on Multi-Objective EAs (MOEAs)
MOOP involves optimizing two or more objectives; they usually conflict with each other (i.e., optimizing one might worsen the second), and it is not possible to retrieve a single solution that fully optimizes all the objectives simultaneously. Therefore, optimality in MOOP is defined by the concept of pairwise dominance: a solution $x_{1}$ dominates a solution $x_{2}$ if $x_{1}$ is not worse than $x_{2}$ in all objectives and $x_{1}$ is strictly better than $x_{2}$ in at least one objective [deb2001multi]. When solutions do not dominate each other, they are considered as optimal. This results in a set of trade-off solutions (i.e., Pareto Front), each with different degrees of gain and losses to the objectives.
EAs are great candidates for tackling MOOPs and looking for a set of solutions due to their population-based nature, allowing them to carry on different alternatives at a time [deb2001multi]. Multi-Objectives Evolutionary Algorithms (MOEAs) aim to retrieve an optimal set (named the Pareto set) with the following properties: (i) convergence, or proximity to the true Pareto Front (which is unknown for engineering problems), and (ii) diversity, ensuring that the solutions are well-spaced so that the Pareto Front is uniformly covered. These properties are satisfied by different types of Multi-Objective Survival Strategies by removing weak solutions from the population. MOEAs usually generate a number of offspring larger than the set population size (to promote space exploration); however, this creates a surplus in the population. Survival strategies deal with this surplus by promoting non-dominated isolated solutions (elites) and removing the ones located in crowded regions to promote diversity. For this study, we will explore two of the most well-known strategies in the state-of-the-art.
Non-Dominated Sorting with Crowding Distance (NS)
Introduced with NSGA-II [deb2002fast], NS (i) divides the population into different sorted non-dominated sets (i.e., the first set dominates all solutions, the second dominates all solutions except the first set, and so on); and (ii) deals with the surplus in the population with the crowding distance operator, which sums the normalized distances between neighboring solutions in each objective dimension it identifies isolated solutions (i.e., larger distances indicate more isolated solutions).
1
Input : Population size n, number of generations g
Output : The non-dominated set of solutions P
2
3begin
4
[…]